# CANVAS: Commonsense-Aware Navigation System for Intuitive Human-Robot Interaction

### v1

@@ I Introduction
Our results show that CANVAS consistently outperforms ROS NavStack [20] across all environments with noisy sketch instructions. Particularly in the challenging orchard environment, CANVAS navigates effectively, while ROS NavStack fails due to its reliance on rule-based algorithms[21]. CANVAS ’s trajectory closely mirrors human demonstrations with fewer commonsense constraint violations, indicating a better understanding of human expectations. Despite being trained only on simulated data, CANVAS also excels in real-world scenarios, demonstrating strong Sim2Real transfer capabilities.
@@ III Dataset and Task
To address this challenge, we introduce COMMAND—a comprehensive experiment suite designed to assess whether robots can use commonsense understanding to transform abstract or occasionally noisy human instructions into the most desired trajectory. COMMAND was collected in three distinct environments: office, street, and orchard, using NVIDIA Isaac Sim [41]. The data consists of high-quality sketch instruction labels and teleoperation data, all by human experts. COMMAND contains 48 hours of driving data covering a distance of 219 kilometers. In this section, we describe the dataset curation process and the task definition that builds upon it.
@@ III-A Dataset
A datapoint in COMMAND includes a canvas map, sketch and language instructions, commonsense constraints, and teleoperation records. The canvas maps, linked to each environment, not only provide occupancy information but also serve as a communication interface between humans and robots. We assume humans provide instructions in two modalities: sketch instructions $S$ that consist of hand-drawn trajectories on maps for the robot to follow, and language instructions $L$ that outline goals and related requirements. When collecting sketch instructions $S$, we introduce both Precise and Misleading conditions to gather training data that enables our model to handle noisy sketch instructions more robustly. In addition to the instructions, we define a set of commonsense constraints $C$, which are derived from the navigation environment $E$ and the language instructions $L$. These constraints help evaluate whether the robot exhibits appropriate navigation behaviors, such as using crosswalks. We also include human teleoperation records to optimize and evaluate the robot behavior against human actions. An overview of our data collection pipeline is illustrated in Figure 2, and the statistics are provided in Table II. We detail each step below.
@@ III-A4 Teleoperation
The data workers are then provided both the sketch and language instructions to teleoperate the virtual robot in NVIDIA Isaac Sim. We record front view images and canvas map images along with the human-drawn trajectory. After collecting the teleoperation data, we adopt the Frechet distance (FD) [43] to measure the discrepancy between the sketch trajectory and the human-demonstrated trajectory. This metric, which indicates noise in the sketch instructions, tends to be higher in the Precise condition and lower in the Misleading condition.
@@ III-B Task Definition
At the end of each iteration, the hindsight trajectory is updated by appending $p(t)$, which represents the robot’s position updated through predicted waypoints $y(t)$, resulting in $H(t+1)=(p(1),p(2),...,p(t))$. This update is reflected in the next canvas map image $X_{c}(t+1)$, while the front view image $X_{f}(t+1)$ is also updated based on the robot’s new position. This process continues until the robot either reaches the destination (within 0.5 meters) or a maximum timestep $t=T$ is reached.
@@ V Experiments
We compare CANVAS with ROS NavStack [20], a straightforward yet effective rule-based navigation system. For this system, we converted the sketch instructions into step-by-step, point-to-point inputs, but language instructions could not be accommodated. The same hyperparameters were used for all experiments with NavStack. We evaluate two variations of CANVAS. CANVAS-S modifies the original Idefics2 8B [45] by swapping the vision encoder for SigLIP-L [46] and the text encoder for Qwen2-0.5B [47], reducing the model size from 8B to 0.7B to better accommodate real-world deployment. In contrast, CANVAS-L retains the original Idefics2 8B [45] architecture with its pre-trained weights. Both models utilize 128 waypoint tokens. In the main experiments, both models were inferred using single NVIDIA H100 GPU. All experiments were evaluated over three iterations for each test dataset, with randomized starting orientations. In the real-world environment, SLAM with FAST-LIO2 [48] was used to find the robot’s current position.
@@ V-A Results in the Simulated Environments
Table  IV compares the IVR between CANVAS and NavStack. In the street (road), commonsense constraints include lane adherence, while in the street (sidewalk), they involve crossing roads correctly and staying on the sidewalk. CANVAS consistently achieves lower IVR, even with misleading instructions, by learning commonsense driving rules from demonstrations, unlike NavStack’s reliance on explicit programming.
@@ V-D Additional Case Study
We perform a qualitative analysis to examine the factors behind the success of CANVAS in comparison to NavStack [20]. Figure 6 highlights typical failure cases for NavStack, where the robot is unable to reach its destination. In 56% of these cases, failures result from stumbling over rocks in the orchard environment. Figure 4 compares CANVAS and NavStack in this primary failure scenario. The orchard has uneven terrain, and NavStack struggles to avoid small but hazardous obstacles like rocks because its limited perception can’t distinguish between rocks and passable areas such as grass. In contrast, CANVAS utilizes visual inputs from the camera to reliably detect unexpected obstacles and assesses their navigation risk based on learned experiences from demonstrations.
Additionally, we assess CANVAS-S and CANVAS-L on their adherence to commonsense constraints. As shown in Figure 4, CANVAS-S disregards the use of a crosswalk, whereas CANVAS-L successfully adheres to the implicit commonsense rule of crossing at designated points.
@@ V-E Real-Time Navigation Feasibility Study
Finally, we evaluate the feasibility of deploying CANVAS in real-time applications. Figure 5 details the robot design. CANVAS demonstrates real-time inference capabilities, with an average latency of 400ms for CANVAS-S and 800ms for CANVAS-L, all within the available memory limits. These results highlight CANVAS’s potential to efficiently handle real-world navigation tasks without substantial delays.

### v2

@@ I Introduction
Our results show that CANVAS consistently outperforms ROS NavStack [20] across all environments with noisy sketch instructions. Particularly in the challenging orchard environment, CANVAS navigates effectively, while NavStack fails due to its reliance on rule-based algorithms[21]. CANVAS ’s trajectory closely mirrors human demonstrations with fewer commonsense constraint violations, indicating a better understanding of human expectations. Despite being trained only on simulated data, CANVAS also excels in real-world scenarios, demonstrating strong Sim2Real transfer capabilities.
@@ III Dataset and Task
To address this challenge, we introduce COMMAND—a comprehensive experiment suite designed to assess whether robots can use commonsense understanding to transform noisy or abstract human instructions into the most desired trajectory. COMMAND was collected in three distinct environments: office, street, and orchard, using NVIDIA Isaac Sim [41]. The data consists of high-quality sketch instruction labels and teleoperation data, all by human experts. COMMAND contains 48 hours of driving data covering a distance of 219 kilometers. In this section, we describe the dataset curation process and its resulting task definition.
@@ III-A Dataset
A datapoint in COMMAND includes a canvas map, sketch and language instructions, commonsense constraints, and teleoperation records. The canvas maps, linked to each environment, provide occupancy information and serve as a human-robot communication interface. We assume humans provide instructions in two modalities: sketch instructions $S$ that consist of hand-drawn trajectories on maps for the robot to follow, and language instructions $L$ that outline goals and related requirements. When collecting sketch instructions $S$, we introduce both Precise and Misleading conditions to gather training data that enables our model to handle noisy sketch instructions more robustly. In addition to the instructions, we define a set of commonsense constraints $C$, which are derived from the navigation environment $E$ and the language instructions $L$. These constraints help evaluate whether the robot exhibits appropriate navigation behaviors, such as using crosswalks. We also include human teleoperation records to optimize and evaluate the robot behavior against human actions. An overview of our data collection pipeline is illustrated in Figure 2, and the statistics are provided in Table II. We detail each step below.
@@ III-A4 Teleoperation
The data workers are then provided both the sketch and language instructions to teleoperate the virtual robot in NVIDIA Isaac Sim. We record front view images, canvas map images, and the human-demonstrated trajectory. After collecting the teleoperation data, we adopt the Frechet distance (FD) [43] to measure the discrepancy between the sketch trajectory and the human-demonstrated trajectory. This metric, which indicates noise in the sketch instructions, tends to be higher in the Precise condition and lower in the Misleading condition.
@@ III-B Task Definition
At the end of each iteration, the hindsight trajectory is updated by appending $p(t)$, which represents the robot’s position updated through predicted waypoints $y(t)$, resulting in $H(t+1)=(p(1),p(2),...,p(t))$. This update is reflected in the next canvas map image $X_{c}(t+1)$, while the front view image $X_{f}(t+1)$ is also updated based on the robot’s new position. This process continues until the robot either reaches the destination or a maximum timestep $t=T$ is reached.
@@ V Experiments
We compare CANVAS with ROS NavStack [20], a straightforward yet effective rule-based navigation system. For this system, we converted the sketch instructions into step-by-step, point-to-point inputs, but language instructions could not be accommodated. The same hyperparameters were used for all experiments with NavStack. We evaluate two variations of CANVAS. CANVAS-S modifies the original Idefics2 8B [45] by swapping the vision encoder for SigLIP-L [46] and the text encoder for Qwen2-0.5B [47], reducing the model size from 8B to 0.7B to better accommodate real-world deployment. In contrast, CANVAS-L retains the original Idefics2 8B [45] architecture with its pre-trained weights. Both models were trained using LoRA [48] with $r$=256, $a$=512, and a dropout rate of 0.1. We use the AdamW optimizer [49] with a different learning rates of 2e-5 for the LLM and 5e-5 for the projector and vision encoder, a batch size of 32, and 5 training epochs. Each model utilizes 128 waypoint tokens. In the main experiments, CANVAS models were inferred using a single NVIDIA H100 GPU. All experiments were evaluated over three iterations for each test dataset, with randomized starting orientations. In the real-world environment, SLAM with FAST-LIO2 [50] was used to find the robot’s current position.
@@ V-A Results in the Simulated Environments
Table  IV compares the IVR between CANVAS and NavStack. In the street (road), commonsense constraints include lane adherence, while in the street (sidewalk), they involve crossing roads correctly and staying on the sidewalk. Unlike NavStack’s rule-based approach, CANVAS learns commonsense driving rules from human demonstrations, consistently achieves lower IVR even with misleading instructions.
@@ V-D Additional Case Study
We perform a qualitative analysis to examine the factors behind the success of CANVAS compared to NavStack [20]. Figure 6 highlights typical failure cases for NavStack, where the robot is unable to reach its destination. In 56% of these cases, failures result from stumbling over rocks in the orchard environment. Figure 4 compares CANVAS and NavStack in this primary failure scenario. The orchard has uneven terrain, and NavStack struggles to avoid small but hazardous obstacles like rocks because its limited perception cannot distinguish rocks from passable areas such as grass.
Tuning the hyperparameters of NavStack exacerbates the issue: setting the obstacle’s height threshold lower causes the system to misclassify grass as an obstacle, and the robot LiDAR sensor, which suffers from significant vibration, struggles to navigate with a small threshold. In contrast, CANVAS utilizes visual inputs from the camera to reliably detect unexpected obstacles and assesses their navigation risk based on learned experiences from demonstrations.
@@ V-E Real-Time Navigation Feasibility Study
Finally, we evaluate the feasibility of deploying CANVAS in real-time applications. The NVIDIA AGX Orin, integrated into our robot, demonstrates on-device inference capabilities with an average latency of 400ms for CANVAS-S and 800ms for CANVAS-L, highlighting its potential for efficient real-world navigation tasks without substantial delays.